OpenStudy (anonymous):

Model the function rule y = -2x - 3 with a table of values.

jimthompson5910 (jim_thompson5910):

pick any value you want for x x = 0 is a good start y = -2x - 3 y = -2(0) - 3 ... plug in x = 0 y = 0 - 3 y = -3

jimthompson5910 (jim_thompson5910):

so when x =0, y is y = -3

jimthompson5910 (jim_thompson5910):

that's one entry in your table

jimthompson5910 (jim_thompson5910):

repeat this for various x values to find other rows in the table